OpenStudy (anonymous):

Rationalize the denominator. assume that all variables represent positive real. (4)/(√(15))

sam (.sam.):

\[\frac{4\sqrt{15}}{15}\]

OpenStudy (anonymous):

2 radical 15

OpenStudy (anonymous):

nvm sorry

OpenStudy (anonymous):

Would you give me an explanation? Thx!!

OpenStudy (anonymous):

I am sorry, I meant Could you??

OpenStudy (saifoo.khan):

\[\frac{4 } {\sqrt{15}} \times \frac{\sqrt{15}}{\sqrt{15}}\] Solve.

OpenStudy (anonymous):